Multivariable Calculus (Math 214) Assignment

due January 26, 2001

1. Let A and B be the points in Cartesian 3-space given by

A = (2, −3, 5) , B = (0, 1, −1) .

(a) Find the sum A + B.

(b) Find the difference A − B.

(c) Find the norms ||A|| and ||B||.

(d) Find the distance of A from B.

(e) What is the angle at the origin in the triangle with vertices at A, B, and the origin?

2. Let Δ be the triangle in the Cartesian plane with vertices

A = (2, 1), B = (1, 3), C = (−1, 2).

Find the point where the 3 medians of Δ intersect.

Document network location:

DVI: http://math.albany.edu:8000/math/pers/hammond/course/mat214/assgt/010123.dvi
HTML: http://math.albany.edu:8000/math/pers/hammond/course/mat214/assgt/010123.html

[Processed from GELLMU to L^{A}TEX on Tue 23 Jan 2001 at 12:14:51 EST]